# Degrees of interior polynomials and parking function enumerators

### v2

@@ 5. Parking function enumerators and greedoid polynomials
In [20] it is proved that the parking function enumerator of an Eulerian digraph can be expressed as the interior polynomial of the cographic matroid. (Prior to that, [13] settled the planar case, which in turn extended [11, Corollary 5.9].) Hence the results of the previous section give us information on the degree of the parking function enumerator of an Eulerian digraph and, equivalently, on the degree of the lowest term of the greedoid polynomial. It turns out that similar results hold for all directed graphs.
@@ 5.1. Preliminaries on greedoids and parking functions
[6, Theorem 1.3] $\lambda_{G,s}(x)=x^{|E|-|V|+1}{\rm park}_{G,s}(x^{-1})$
@@ 5.3. Arbitrary directed graphs
Swee Hong Chan points out that in the case of directed branching greedoids, the greedoid activity notion of [5] specializes to the following.
@@
Recall that ${\rm minfas}(G,s)$ is the minimal cardinality of an edge set whose removal leaves an $s$-root-connected acyclic digraph (Definition 1.9). It follows from Theorems 1.8 and 1.10 that for a connected Eulerian digraph and an arbitrary vertex $s$, ${\rm minfas}(G)={\rm minfas}(G,s)$. It is also not very hard to prove this directly.

### v3

@@ 5. Parking function enumerators and greedoid polynomials
In [22] it is proved that the parking function enumerator of an Eulerian digraph can be expressed as the interior polynomial of the cographic matroid. (Prior to that, [15] settled the planar case, which in turn extended [13, Corollary 5.9].) Hence the results of the previous section give us information on the degree of the parking function enumerator of an Eulerian digraph and, equivalently, on the degree of the lowest term of the greedoid polynomial. It turns out that similar results hold for all directed graphs.
@@ 5.1. Preliminaries on greedoids and parking functions
[7, Theorem 1.3] $\lambda_{G,s}(x)=x^{|E|-|V|+1}{\rm park}_{G,s}(x^{-1})$
@@ 5.3. Arbitrary directed graphs
Swee Hong Chan [7] points out that in the case of directed branching greedoids, the notion of greedoid activity can be rephrased in a more graph theoretic way. For this we need some additional definitions.
Let $T$ be a spanning tree in the digraph $G$. For an edge $e\notin T$, we say that $e^{\prime}\in C(T,e)$ stands parallel to $e$ in the fundamental cycle $C(T,e)$ if $e$ and $e^{\prime}$ point in the same cyclic direction. Otherwise we say that $e^{\prime}$ stands opposite to $e$. (See Example 5.16.)
@@
Recall that ${\rm minfas}(G,s)$ is the minimal cardinality of an edge set whose removal leaves an $s$-root-connected acyclic digraph (Definition 1.10). It follows from Theorems 1.9 and 1.11 that for a connected Eulerian digraph and an arbitrary vertex $s$, ${\rm minfas}(G)={\rm minfas}(G,s)$. It is also not very hard to prove this directly.